OpenStudy (anonymous):

what are the real zeros of x^2+12x+36

OpenStudy (anonymous):

x=-6 and here some details http://www.wolframalpha.com/input/?i=x%5E2%2B12x%2B36&t=crmtb01

OpenStudy (anonymous):

\[x^2+12x+36 = (x+6)(x+6)\] First you factor, then set the factored expression equal to zero. \[(x+6)(x+6) = 0\] Solve for x. Because anything multiplied by zero is zero, set each factor equal to zero and solve for x. Because both factors are the same, you just need to do this once. \[x+6 =0\\x=-6\]

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

This becomes useful later when you are trying to find maximum and minimum values on curves.

OpenStudy (anonymous):

move your x to one side positive become negative

OpenStudy (anonymous):

Are you following, @rhines8 ?

OpenStudy (anonymous):

yes. Thank You!

OpenStudy (anonymous):

:D Have a nice day.